Sorry, you need to enable JavaScript to visit this website.

MCST Banner Image

am1weekly_2004-297_2004-303

Mission Operations Days: 2004/297 to 2004/303
October 23, 2004 00:00:00 GMT to October 29, 2004 00:00:00 GMT


Terra Spacecraft and MODIS Instrument Status:

AM-1 "Terra" is in Normal Mode
PFM MODIS is in A-Side Nominal Science Mode
(Cross Strapped with FR B)
ComponentState(s)Comment
BlackbodyA On (290K); B OffNominal Science Mode
Calibration ElectronicsA On; B OffNominal Science Mode
Control ProcessorA On; B OffNominal Science Mode
Nadir DoorUnlatched, OpenNominal Science Mode
Space View DoorUnlatched, OpenNominal Science Mode
Solar Diffuser DoorUnlatched, ScreenedPermanent Configuration
FDDI FormatterA On; B OffNominal Science Mode
FDDI PortA On; B OffNominal Science Mode
FIFO Memory#1 & #2 On; #3 & #4 OffNominal Science Mode
Format ProcessorA Off; B OnNominal Science Mode
Power Supplies#1 On, #2 OffNominal Science Mode
PV VIS FPAA On; B OffNominal Science Mode
PV NIR FPAA On; B OffNominal Science Mode
PV SMIR FPAA On; B OffNominal Science Mode
PV LWIR FPAA On; B OffNominal Science Mode
PC LWIR FPAA On; B OffNominal Science Mode
RC Outgas HeatersAll 3 (CS, IS, OS) OffNominal Science Mode
RC LWIR FPA HeaterOn (83K)Nominal Science Mode
RC SMIR FPA HeaterOffNominal Science Mode
Scan AssemblyA On; B OffNominal Science Mode
SDSMOffNominal Science Mode
SRCAOffNominal Science Mode
PS1 Survival HeaterEnabledNominal Science Mode
PS2 Survival HeaterEnabledNominal Science Mode
Timing GeneratorA On; B OffNominal Science Mode
Flight SoftwareRev BD + 11 sets of patches
  • CP07 to Survival Mode - No SDD to closed command (1 patch)
  • CP07 to Safe Mode - No FR OFF Command (1 patch)
  • CP07 to Safe Mode - No SDD to closed command (1 patch)
  • CP07 to Science Mode - Turn on FRB instead of FRA (1 patch)
  • CP07 to Outgas and Science Mode - Turn on FRB instead of FRA (1 patch)
  • FR Error Counting and Statistics (2 patches)
  • FR 1000ms time code patches (3 patches)
  • FR Band 5 Gain Change (20 commands)
  • FR BB Ideal Output Values (1 patch)
  • FR A EEPROM KME Fast 1 word patch (1 patch)
  • FR B EEPROM KME Fast 1 word patch (1 patch)
Nominal Science Mode
Inhibit Ids SetnoneNominal Science Mode
TMON 66DisabledNominal Science Mode
TMON 67DisabledNominal Science Mode
TMON 68EnabledNominal Science Mode
TMON 69DisabledNominal Science Mode

This Week’s Completed MODIS Activities:

Saturday, October 23, 2004

2004/297 21:00:00 ATC MOD_OA26_BB_270K_A.1 Set blackbody temperature to 270K on the A Side

Sunday, October 24, 2004

2004/298 15:00:00 ATC MOD_OA26_BB_280K_A.1 Set blackbody temperature to 280K on the A Side
2004/298 19:00:00 ATC MOD_OA26_BB_285K_A.1 Set blackbody temperature to 285K on the A Side
2004/298 22:00:00 ATC MOD_OA26_BB_290K_A.1 Set blackbody temperature to 290K on the A Side

Monday, October 25, 2004

2004/299 02:00:00 ATC MOD_OA26_BB_315K_A.1 Set blackbody temperature to 315K on the A Side
2004/299 02:00:30 ATC MOD_OA26_BB_DCYCLE_FULL.1 Set blackbody heater duty cycle to FULL
2004/299 05:00:00 ATC MOD_OA26_BB_270K_A.1 Set blackbody temperature to 270K on the A Side
2004/299 05:00:30 ATC MOD_OA26_BB_DCYCLE_THIRD.1 Set blackbody heater duty cycle to THIRD

Tuesday, October 26, 2004

2004/300 02:00:00 ATC MOD_OA26_BB_290K_A.1 Set blackbody temperature to 290K on the A Side
2004/300 20:09:32-20:18:20 ATC MOD_OA16_SDSM_SCRN_NO_DOOR_A.4 OA-16: SD/SDSM screened side A, no door movement.
2004/300 20:50:00-20:50:31 ATC MOD_SRCA_11_3_8_2_SETUP.1 SRCA setup activity using lamp #4 in lieu of lamp #2
2004/300 20:58:49-21:19:44 ATC MOD_OA20_SRCA_10W_RAD_A_Base.2 OA-20: SRCA 10W Continuous Radiometric - A Side, Using Lamp #1
Data Collect from 21:14:43-21:19:44

Wednesday, October 27, 2004

2004/301 03:20:00-03:20:31 ATC MOD_SRCA_NOMINAL_SETUP.4 Sets up nominal A Side SRCA lamp configuration
2004/301 03:31:40-04:23:21 ATC MOD_OA22_SRCA_FULL_SPECT_1A.6 OA-22: SRCA Full Spectral Calibration Part 1 - A Side
30W lamp data collection is from 03:47:34-04:23:20
2004/301 05:27:25-06:10:15 ATC MOD_OA22_SRCA_FULL_SPECT_2A.5 OA-22: SRCA Full Spectral Calibration Part 2 - A Side
30W lamp data collection is from 05:27:25-06:01:55
2004/301 06:40:00-06:40:31 ATC MOD_SRCA_NOMINAL_SETUP.4 Sets up nominal A Side SRCA lamp configuration
2004/301 06:49:27-07:40:45 ATC MOD_OA22_SRCA_FULL_SPECT_3A.6 OA-22: SRCA Full Spectral Calibration Part 3 - A Side
10W lamp data collection is from 07:05:54-07:40:27
2004/301 08:41:28-09:26:45 ATC MOD_OA22_SRCA_FULL_SPECT_4A.4 OA-22: SRCA Full Spectral Calibration Part 4 - A Side
10W lamp data collection is from 08:41:28-09:19:28

Thursday, October 28, 2004

None

Friday, October 29, 2004

None

This Week’s Scheduled MODIS Activities Not Completed:

Tuesday, October 26, 2004

For Terra inclination adjust maneuver #10:
2004/300 15:11:29-15:13:10 ATC MOD_CLOSE_NAD_SVD.1
(100 seconds prior to S/C day, orbit 25838)

Thursday, October 28, 2004

For Terra inclination adjust maneuver #11:
2004/302 14:59:17-15:00:58 ATC MOD_CLOSE_NAD_SVD.1
(100 seconds prior to S/C day, orbit 25867)

Upcoming MODIS Events:

Calibrations for November will include biweekly SDSM, monthly SRCA Full and 10W Radiometric, and bimonthly SRCA Full Spatial calibrations.

A request by MCST has been made to perform SRCA 10W and 1W continuous mode Radiometric observations. These activities will be coordinated for November. These tests will be similar to the ones perform in October 2003.

Tuesday, November 2, 2004

Lunar Roll Maneuver (-14.1615 deg roll)
2004/306 23:44:52 ATC MOD_OA03_FR_NIGHT_RATE (03:23 early)
2004/307 00:22:14-00:29:00 ATC MOD_OA08_ROLL_MAN
2004/307 00:41:04-01:27:09 ATC MOD_OA03_FR_DAY_RATE (03:23 late)

Thursday, November 4, 2004

2004/309 15:06:08-15:07:49 ATC MOD_CLOSE_NAD_SVD.1
(100 seconds prior to S/C day, orbit 25968)

Friday, November 5, 2004

2004/310 01:45:00-01:45:31 ATC MOD_SRCA_14_6_4_1_SETUP.1 Turns on the SRCA and sets it's lamp config. to use 10W lamp #2
2004/310 01:50:09-03:54:57 ATC MOD_OA20_SRCA_10W_RAD_A_Base.2 Performs SRCA 10W continuous radiometric activity
Data collect from 02:06:03-03:54:57
2004/310 08:35:00-08:35:31 ATC MOD_SRCA_14_6_4_1_SETUP.1 Turns on the SRCA and sets it's lamp config. to use 10W lamp #2
2004/310 08:40:06-10:31:32 ATC MOD_SRCA_SPECIAL_10WND.1 Performs SRCA 10W continuous radiometric activity with the neutral density filter in place
Data collect from 08:41:10-10:31:32
2004/310 14:55:00-14:55:31 ATC MOD_SRCA_14_6_4_1_SETUP.1 Turns on the SRCA and sets it’s lamp config. to use 1W lamp #1 (backup)
2004/310 15:01:16-17:06:04 ATC MOD_OA21_SRCA_1W_RAD_A.5 Performs SRCA 1W continuous radiometric activity
Data collect from 15:17:10-17:06:04

Maneuvers:

The November Lunar Calibration Roll will be on day 2004/307 (November 2, 2004). Data collect will be from 00:22:22- 00:28:52 GMT with a roll angle of -14.1615 degrees.

The Terra FOT has successfully performed 2 Inclination Adjust Maneuvers (IAMs), 2004/300 and 2004/302, respectively. A third IAM is planned for 2004/309 and will occur at approximately 15:00 GMT.

MODIS Anomalies:

MODIS NAD (Nadir Aperture Door) Anomaly:

Telemetry still indicates that the “NAD Open” (MOD_CR_DR_NAD_OPEN) trigger is set to the “NOT_OPEN” position; however, science data has shown that the NAD is open. The NAD OPEN trigger may yet “pop” to OPEN due to thermal cycling on the trigger, as was seen on 2002/079.

General Instrument Comments:

MODIS is in Science Mode, using all A-side components, but cross-strapped to Formatter B, with the blackbody maintained at 290K and the Solar Diffuser Door at the Screened position.

MODIS Telemetry Trends:

Nominal telemetry trending is ongoing.

Non-MODIS Significant Events:

Terra Solid State Recorder Anomaly

Overview:

The SSR Anomaly is still under investigation. MODIS currently is operating with 34 Supersets allotted for its buffer. There are plans in the works for possibly power cycling the SSR to get back PWA’s that are in the Bad List, but no date has yet been set. There are no current plans for a change in science data capture rates. For details, see previous report am1weekly_2003-305_2003-311.doc. The Terra FOT is currently evaluating plans to perform a power recycling of the SSR in an attempt to regain functionality of the “bad” PWAs. This activity will be performed at a time when it is least detrimental to critical science data.

The EMOS mission status webpage (which includes instrument data loss records) can be found at the following URL. Note that a password is now required to access the EOS mission links web pages. There are directions at the URL below for how to obtain a username/password.

http://jupiter02.gsfc.nasa.gov/links/links.htm

Limited Life Item Status:

SRCA 10W Lamp #1: 237.4 of 500 hours
SRCA 10W Lamp #2: 168.7 of 500 hours
SRCA 10W Lamp #3: 182.1 of 500 hours
SRCA 10W Lamp #4: 67.0 of 500 hours

SRCA 1W Lamp #1: 566.9 of 4000 hours
SRCA 1W Lamp #2: 276.9 of 4000 hours

Solar Diffuser Door: 2146 of 3022 Movements
Nadir Aperture Door: 538 of 1316 Movements
Space View Door: 441 of 1316 Movements